OpenStudy (anonymous):

SIMPLIFY? -9√28a^2 * 1/3√63a

OpenStudy (anonymous):

@ganeshie8 could you help?

OpenStudy (anonymous):

\[-9\sqrt{28a^2}\times \frac{1}{3}\sqrt{63a}\]?

OpenStudy (anonymous):

@satellite73 yes :)

OpenStudy (anonymous):

k \[28=4\times 7\] so \[\sqrt{28a^2}=\sqrt{4a^2}\sqrt7=2a\sqrt7\] for the first term

OpenStudy (anonymous):

\[63=9\times 7\] so \[\sqrt{63a}=\sqrt9\sqrt{7a}=3\sqrt{7a}\] for the second term then we can multiply

OpenStudy (anonymous):

\[-9\times 2a\sqrt7\times \frac{1}{3}\times 3\sqrt{7a}\] \[=-18a\sqrt{49a}=-18\times 7\times \sqrt{a}=-126\sqrt{a}\]

OpenStudy (anonymous):

@satellite73 okay that helped so much! thank you :)

OpenStudy (anonymous):

yw